# Geometric Self-Supervised Pretraining on 3D Protein Structures using Subgraphs

### v2

@@ 2 Related Work
Protein representation learning aims to learn informative embeddings that capture the biological and functional characteristics of proteins. Early methods primarily focused on sequence-based representations [34, 35]. Recent advancements have shifted towards multimodality, by integrating the structural information of proteins. For instance, methods like HoloProt [36] incorporate sequence, surface and structure information, DeepFRI [10] propose a GCN to solve protein function prediction tasks while GAT-GO [37] introduces an attention-based graph model. Moreover, with the advance of language models, the researches have started integrating and encoding also text information for the proteins such as Prot2Text [15], ProtST [38] and ProteinDT [39]. 3D GNNs have also emerged as a promising approach to capture the spatial relationships within protein structures. wang2023learning [12] introduced ProNet, a 3D GNN model that integrates spatial and geometric information for protein classification tasks. schutt2018schnet [31] developed SchNet, which incorporates radial basis functions to handle pairwise distances in molecular graphs. coors2018spherenet [32] proposed SphereNet, a spherical representation of molecular structures that enhances spatial encoding. Our work is orthogonal to these methods, as it can be applied to various backbone architecture, aiming to improve the learned representations by leveraging the geometric structure in 3D protein data throughr pretraining.
@@ 3.2 Geometric Self-Supervised Pretraining
$\begin{split}{\bm{c}}_{G}&=\frac{1}{|V|}\sum_{i\in V}{\bm{p}}_{i}\\ {\bm{c}}_{G}&=\left(\frac{1}{|V|}\sum_{i\in V}x_{i},\frac{1}{N}\sum_{i\in V}y_{i},\frac{1}{|V|}\sum_{i\in V}z_{i}\right)\end{split}$ (5)

### v3

@@ 2 Related Work
Protein representation learning aims to learn informative embeddings that capture the biological and functional characteristics of proteins. Early methods primarily focused on sequence-based representations [20, 23]. Recent advancements have shifted towards multimodality, by integrating the structural information of proteins. For instance, methods like HoloProt [32] incorporate sequence, surface and structure information, DeepFRI [9] propose a GCN to solve protein function prediction tasks while GAT-GO [21] introduces an attention-based graph model. Moreover, with the advance of language models, the researches have started integrating and encoding also text information for the proteins such as Prot2Text [1], ProtST [44] and ProteinDT [22]. 3D GNNs have also emerged as a promising approach to capture the spatial relationships within protein structures. Wang et al. [40] introduced ProNet, a 3D GNN model that integrates spatial and geometric information for protein classification tasks. Schütt et al. [30] developed SchNet, which incorporates radial basis functions to handle pairwise distances in molecular graphs. Coors et al. [4] proposed SphereNet, a spherical representation of molecular structures that enhances spatial encoding. Our work is orthogonal to these methods, as it can be applied to various backbone architecture, aiming to improve the learned representations by leveraging the geometric structure in 3D protein data throughr pretraining.
@@ 3.2 Geometric Self-Supervised Pretraining
$\begin{split}{\bm{c}}_{G}&=\frac{1}{|V|}\sum_{i\in V}{\bm{p}}_{i}\\ {\bm{c}}_{G}&=\left(\frac{1}{|V|}\sum_{i\in V}x_{i},\frac{1}{|V|}\sum_{i\in V}y_{i},\frac{1}{|V|}\sum_{i\in V}z_{i}\right)\end{split}$ (5)